Circular No. 8554

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 JD_{108} (CATALINA-NEAT) K. J. Lawrence, Jet Propulsion Laboratory, reports the NEAT discov-

ery of a comet on NEAT images taken on June 28.4 UT, the object showing a fan-shaped tail ∼ 10′′ long in p.a. 270^{o}. Following posting on the “NEO Confirmation Page”, J. E. McGaha (Tucson, AZ, 0.62-m f/5.1 reflector) reports that his 300-s CCD exposures taken on June 30.33–30.35 show a stellar nucleus with a uniform round coma 7′′ in diameter and a 4′′-long fan-shaped tail in p.a. 215^{o}. J. Young (Table Mountain, 0.6-m reflector) adds that his CCD frames taken on June 30.4 show a small but bright 5′′- diameter coma and a faint, straight 16′′ tail in p.a. 255^{o}. The Minor Planet Center has noted that this object was observed incidentally (as apparently asteroidal) by the Catalina Sky Survey on May 12, by LONEOS on May 13, and by the NEAT survey on June 21 — the first two nights of observations being published with the designation 2005 JD_{108} on MPS 132975.

2005 UT α_{2000} δ_{2000} Mag. Observer
May 12.42963 21^{h}10^{m}23.^{s}71 −12^{o}45′21′′.7 18.9 Catalina
June 28.39944 21 16 54.32 −11 31 20.6 18.4 NEAT

The new astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2005-M65.

T = 2005 Aug. 1.3467 TT  ω = 88.^{o}9493
e = 0.376713  Ω = 224.3811 }2000.0
q = 4.032734 AU  i = 3.2761
a = 6.470112 AU n^{o} = 0.0598875 P = 16.46 years

THE EDGAR WILSON AWARD 2005 The Smithsonian Astrophysical Observatory announces that the 2005

Edgar Wilson Award for the discovery of comets (cf. IAUC 6936, 8372) is being divided among the following two individuals: Roy A. Tucker, Tucson, AZ, U.S.A., for C/2004 Q1; and Donald Edward Machholz, Jr., Colfax, CA, U.S.A., for C/2004 Q2.

COMET P/2005 JQ_{5} (CATALINA) J. J. Gonzalez, Leon, Spain, reports that he has detected this comet

visually at total mag 10.3–10.5 (coma diameter 4′) on June 25.94 (altitude 11^{o}) and 26.94 UT (altitude 8^{o}) with a 0.20-m reflector.

2005 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green